OpenStudy (anonymous):

MEDAL IF SOLVED CORRECTLY!!!!!!!!!!!! What is the vertex of the equation: y = x^2 +10x +24

OpenStudy (anonymous):

We need to use formula for vertex ( -b/2a) can you try ?

OpenStudy (mathmale):

Rene, it sounds as tho' you're hoping that someone else will solve this for you. That's not a realistic expectation. How would YOU begin the solution of this problem? What have you thought of so far? There's a formula for the x-coordinate of the vertex, by the way; are you familiar with it?

OpenStudy (anonymous):

its an example question but my books instructions are vague, so i need help to understand

OpenStudy (anonymous):

Compare given equation with general quadratic equation y =ax^2+bx+ c So a =----- ? b = --- ? c = ---- ? can you try ?

OpenStudy (anonymous):

so \[1^{2}+10x+24? \]

OpenStudy (anonymous):



OpenStudy (mathmale):

suyogsomavanshi shared the following with us: We need to use formula for vertex ( -b/2a) This is all you need to find the x-coordinate of the vertex.

OpenStudy (mathmale):

the same person asked you to identify the values a, b and c in y = x^2 +10x +24. Can you do that? Hint: the equation could be \[y = 1x^2 +10x +24.\] written as

OpenStudy (mathmale):

a=? b =? c=? x=-b/a = ?

OpenStudy (anonymous):

yes first we need to find out x co ordinate of vertex and then we have to plug it into given equation to find out y co ordinate of vertex

OpenStudy (mathmale):

fearlessrene: Please get involved. Even if you're unsure, tell us what values you think a, b and c have. Or ask questions. But please be heard!

OpenStudy (mathmale):

Live up to your name, fearlessrene. Sometimes we have to take risks.

OpenStudy (mathmale):

Hint: in y = x^2 +10x +24, c = 24. a=? b=?

OpenStudy (loser66):

let me interpret it \[y = \color{red}{a}x^2+\color{blue}{b}x +\color{green}{c}\] your problem is \[y = \color{red}{1}x^2+\color{blue}{10}x +\color{green}{24}\] I think you can answer the question of @mathmale which is "what are a, b, c on your problem?"

OpenStudy (mathmale):

@FearlessRene : Please, would you join this conversation? It's YOUR conversation.